Will Rate:P13.8B (LO 3) (Premium Entries) To stimulate the sales of its Royal Frozen dinners, King Company places 1 coupon in each box.
Answer #1

Ans:

The following will be the journal entries to be recorded relative to premium plan.

S.no

Particulars

Debit

Credit

1.

Premium inventory a/c

225,000

       Cash a/c

225,000

(being hot meal carriers being purchased 100,000 meal carriers @$2.25)

2.

Cash a/c

1,04,50,000

          sales revenue a/c

1,04,50,000

(being sales of royal frozen dinners @$4.75 a dinner is recorded)

3.

Premium expense a/c

51,750

           Premium inventory

51,750

(being expense for 230,000 coupons presented calculated as (230,000 / 10)*$2.25 = $51,750)

4.

Premium expense a/c

$96,750

            Premium liability

$96,750

(being unexpired liability on premium expense recorded, it is calculated as (22,00,000*30% = 660,000 - 230,000 = 430,000, (430,000/10)*$2.25 = 96,750)